MDRS Daily Summary Report for Sol 01
Summary Title Crew 162 Team ISU is on Mars!
Mission Status: Sol 01 went well, crew is in good spirits and worked
hard today on set-up so that the rest of the mission will run smoothly.
Overall mission status: good.
Sol Activity Summary: Our team helped Crew 161 pack and depart from the
Hab. We settled into our new home for the next two weeks, performed
inventories and research equipment set-up.
Look Ahead Plan: We are focusing on research project set-up and
initialization this week, filming introduction videos for education and
outreach and putting together or daily work plans to maximize our
productivity while in sim. We have a planned news media visit at the end
of our first week and an international photographer visit during our
Anomalies in work: No problems experienced on Sol 01
Weather: Slightly cloudy and cold morning. High winds experienced in the
afternoon. Dry and cold evening.
Crew Physical Status: Health & Status Officer performed initial overall
health exams on the crew and determined them to be fit for sim. At the
end of the day, dehydration headaches were reported by two crew so we
will focus on hydrating early and often and resting in the evenings.
EVA: No EVA performed on Sol 01.
Reports to be filed:
- Commander’s Report
- Green Hab Report
- EVA Request
- Engineering Report
- HSO Initial Report
- Journalist Report
- Astronomy Report
- Team head shots
- Mission Plan
Support Requested: Nothing requested beyond nominal comms window